OpenStudy (anonymous):

use the elimination method to solve the system of equations 5u+3v = -18 2u + v =-6

jhonyy9 (jhonyy9):

multiply second equation by -3 5u+3v=-18 -6u-3v=18 ------------ -u 0 =0 u=0 5*0 +3v= -18 0 3v= -18 v= -18/3 = -6 u=0 v=-6